1. The electric force between two point charges with a magnitude of 800uC and 900uC is 15 N. How far apart are the two charges in cm?


2. A force of 500 N between two identical point charges is separated by a distance of 47m. Calculate the magnitude of the two point charges.


1
Expert's answer
2022-03-21T14:06:36-0400

1.

"r=\sqrt{\frac{kq_1q_2}{F}}\\
=\sqrt{\frac{9*10^9*800*10^{-6}*900*10^{-6}}{15}}=21\:\rm m"

2.

"q=\sqrt{\frac{Fr^2}{k}}\\
=\sqrt{\frac{500*47^2}{9*10^9}}=0.011\:\rm C"